OpenStudy (anonymous):

prove that the curves y^2 = 4x and x^2 = 4y divide the area of the square bounded by x=0 ,x=4 and y = 4 and y = 0 in three equal parts?

OpenStudy (anonymous):

i think this is from ncert application of integration

OpenStudy (anonymous):

the same sum is done in the book's solved example

OpenStudy (anonymous):

any way area enclosed by the parabola x^2 = 4yand the x-axis is obtained by integrating (x^2/4)dx from limits 0 to 4 its integral is [x^3/12] putting the limits we have [4^3/12] - [0^3/12]=16/3

OpenStudy (anonymous):

similarly area enclosed by the parabola y^2 = 4x and the y-axis is obtained by integrating (y^2/4)dy from limits 0 to 4 its integral is [y^3/12] putting the limits we have [4^3/12] - [0^3/12]=16/3

OpenStudy (anonymous):

hence the third part =4*4 -(16/3) -(16/3) =16/3

OpenStudy (anonymous):

the curves y^2 = 4x and x^2 = 4y divide the area of the square bounded by x=0 ,x=4 and y = 4 and y = 0 in three equal parts and each part equals 16/3

OpenStudy (anonymous):



OpenStudy (anonymous):

now just draw tow parabolas

OpenStudy (anonymous):

by the way x=0 is the eq for y-axis and y=0 is the eq for x-axis

OpenStudy (anonymous):